4. The brick, with mass m = 2 kg is placed on a spring scale. a) Calculate the force of gravity and the weight of the brick. b) How many centimeters does the spring compress? It is known that in order to compress the spring by 1 cm, a force of 5 N must be applied.


1
Expert's answer
2021-02-22T10:26:58-0500

(a) The force of gravity (or the weight of the brick) can be found as follows:


"F_g=W=mg=2\ kg\cdot9.8\ \dfrac{m}{s^2}=19.6\ N."

(b) As we know from the condition of the question, the spring constant equals "5\ \dfrac{N}{cm}" (in order to compress the spring by 1 cm, a force of 5 N must be applied). Therefore, we can find the compression of the spring from the Hooke's Law:


"F_g=mg=kx,""x=\dfrac{F_g}{k}=\dfrac{19.6\ N}{5\ \dfrac{N}{cm}}=3.92\ cm."